OpenStudy (anonymous):

A 28 foot ladder leans up against a wall. How high up in the wall does the ladder reach if the foot is placed 8 feet from the bottom of the wall?

OpenStudy (michele_laino):

here we have to apply the theorem of Pitagora, namely the requested height is given by the subsequent formula: \[\large h = \sqrt {{{28}^2} - {8^2}} = ...?\]

OpenStudy (anonymous):

h=720

OpenStudy (michele_laino):

 what is \[\large \sqrt {720} = ...?\]

OpenStudy (anonymous):

26.832

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!